Two number differ by 2 the sum of their squares is 244 find the number.
Answer by mananth(16949)

let the numbers be x & y
x-y=2
x=(y+2)
x^2+y^2=244
(y+2)^2+y^2=244
y^2+4y+4 +y^2=244
2y^2+4y-240=0
/2
y^2+2y-120=0
y^2+12y-10y-120=0
y(y+12)-10(y+12)=0
(y-10)(y+12)=0
y= 10 OR -12
plug values of y in first equation
x-y=2
when y=10 x=12 (12,10)
when y=-12
x=-10
(-10,-12)